# A Comprehensive Survey of Direct Preference Optimization: Datasets, Theories, Variants, and Applications

### v2

@@ 4 Datasets
Datasets Task Description Dataset Size (#Rows)
@@
Summarize from Human Feedback (Stiennon et al., 2022) Summarization 193,841
OpenAI:Webgpt  (Nakano et al., 2022) Question-answering 19,578
Nvidia/HelpSteer(Wang et al., 2023c) Helpfulness Preference 37,120
Nvidia/HelpSteer2 (Wang et al., 2024b) Helpfulness Preference 21,362
OpenAssistant/oasst1 (Köpf et al., 2023) Assistant Conversations 88,838
Stanfordnlp/SHP (Ethayarajh et al., 2022) Reward-model Training 385,563
Tasksource/Dpo-pairs (Sileo, 2024) Discriminative Tasks 5,128,939
RLHF-V-Dataset (Yu et al., 2024a) Multi-Modal 5,733
HH-RLHF (Bai et al., 2022a) HH preference 169,352
Chatbot Arena (Zheng et al., 2023a) LLM evaluation 33,000
Alpaca Farm Human (Dubois et al., 2024) Instruction-tuning 101,000
@@
RLAIF-V-Dataset (Yu et al., 2024b) Multi-Modal 33,835
Math-Step-DPO (Lai et al., 2024) Mathematical-Reasoning 10,795
UltraFeedback (Cui et al., 2023a) Reward-model Training 63,967
Distilabel-capybara-dpo (Argilla, 2024) Conversation Training 7,563
Chatml-dpo-pairs (Mukherjee et al., 2023) Instruction-Tuning 12,859
Magpie-Llama-3.1-DPO (Xu et al., 2024d) Instruction-Tuning 100,000
Tldr-preference-dpo (davanstrien, 2024) Summarization 522
Zsql-postgres-dpo (Zerolink, 2024a) Text2SQL 259,326
Truthy-dpo (Zerolink, 2024b) Truthfulness Preference 1,016
Py-dpo(Jondurbin, 2024a) Code Generation 9,466
MMInstruction/VLFeedback (Li et al., 2023a) Multimodal 80,258
sqrti/SPA-VL (Zhang et al., 2024a) Multimodal safety 100,788
Gutenberg-DPO (Jondurbin, 2024b) Novel Writing 918
Imdb-dpo (Yuasosnin, 2024) Movie Summarization 10,000
RewardBench (Lambert et al., 2024) Reward model evaluation 8,108
Stack Exchange Preference (Lambert et al., 2023) Instruction-tuning 10,741,532
Nectar (Zhu et al., 2023) Ranking 182,954
OrcaSlimOrca (Lian et al., 2023) Instruction-tuning 517,982
Alpaca Farm GPT-4 (Dubois et al., 2024) Instruction-tuning 20,000
@@
DPO algorithm aligns LLMs using datasets that consist of preference pairs. Typically, such a dataset includes a prompt $x$ paired with two responses: a preferred (chosen) answer and a dispreferred (rejected) answer. Typically, human annotators are guided to evaluate and provide preferences among multiple responses generated from the same prompt. These datasets are denoted as human-labeled datasets in this survey. This process, however, is resource-intensive, especially when dealing with large volumes of data. To mitigate the high cost associated with human annotation, some datasets employ language models to synthesize preference pairs (Cui et al., 2023a) (Yu et al., 2024b) (Zhang et al., 2024a). These synthesized datasets are referred to as AI-labeled datasets in this survey. In this section, we will explore the two types of datasets in greater detail.
@@ 4.1 Human labeled
OpenAI:Summarize From Human Feedback. In downstream task like summarization, public online forum like Reddit can provide abundant text sources with a variety of topics. The TL;DR summarization dataset (Völske et al., 2017) contains 3 millions post from reddit forum. To enhance the quality of data, Stiennon et al. (Stiennon et al., 2022) filtered the dataset and left 123,169 posts with  5% as a validation set. Then  Stiennon et al. (2022) collect human feedback on the filtered TL;DR dataset. To improve human labelers’s responses on the instruction, Stiennon team answered human labeler’s questions in a shared chat room, and provided regular feedback on their performance, which ensures high agreement on the judgements of responses. The new collected human preference dataset contains 64k text summarization examples including human-written responses and human-rated model responses, which could server as high quality sources of training data for DPO.
@@
OpenAI:Webgpt comparisons. Webgpt comparisons (Nakano et al., 2022) consist of 19,578 comparisons that are deemed appropriate for reward modeling, aiming at training a long-form question answering model aligend with human preferences. Each example in the dataset includes a pair of model-generated answers for a given question, along with associated metadata. Each answer is assigned a preference score from human annotators, allowing for the identification of the superior response. To gather these human preferences,  (Nakano et al., 2022) hire well-educated annotators, typically holding at least an undergraduate degree, due to the complexity of the tasks involved. Ultimately, this dataset captures human preferences in the generation of WebGPT, facilitating the optimization of the question-answering model to better align with human expectations.
@@
Nvidia/HelpSteer. HelpSteer (Wang et al., 2023c) is an open-source Helpfulness Dataset (CC-BY-4.0) designed to enhance model alignment for improved helpfulness, factual accuracy, and coherence, while allowing for adjustments in response complexity and verbosity. The dataset comprises 37,120 samples, each containing a prompt, a generated response, and five human-annotated attributes rated from 0 to 4, with higher scores indicating better performance.Prompts are sourced from various origins, with approximately half generated by Scale AI. The remaining prompts come from Open Question Answering, Generation, and Brainstorming tasks, as well as five specific tasks: Rewrite, Summarization, Classification, Extraction, and Closed Question Answering. Responses are produced by an early version of a 43 billion parameter in-house model, which generates up to four responses per prompt using sampling techniques to ensure diversity while maintaining reasonableness. The annotation process involved around 200 U.S.-based human annotators, who evaluated each response based on five attributes: Helpfulness, Correctness, Coherence, Complexity, and Verbosity, using a Likert-5 scale from 0 to 4.
@@
Nvidia/HelpSteer2. HelpSteer2 (Wang et al., 2024b), an open-source helpfulness dataset designed to train state-of-the-art reward models, comprises 21,362 samples, each containing a prompt, a response, and five human-annotated attributes, rated from 0 to 4, with higher scores indicating better performance. Consecutive samples (e.g., sample 1 with sample 2, sample 3 with sample 4) share the same prompt, allowing for preference pairs based on the helpfulness score (e.g., for training DPO or Preference RM), in addition to training SteerLM Regression RM. Over 95% of the prompts are sourced from ShareGPT (RyokoAI, 2023) , a platform where ChatGPT users voluntarily share their conversations, while a small proportion ( 5%) is generated by humans from Scale AI.Responses are produced by early versions of a mix of ten different in-house LLMs, generating two responses per prompt (one from each model) using sampling techniques to provide diverse yet reasonable answers.
@@
OpenAssistant/oasst1. OpenAssistant Conversations (OASST1) (Köpf et al., 2023) is a human-generated, human-annotated assistant-style conversation corpus comprises 161,443 messages in 35 different languages, annotated with 461,292 quality ratings, resulting in over 10,000 fully annotated conversation trees. The corpus is the product of a worldwide crowd-sourcing initiative involving over 13,500 volunteers.The dataset consists of message trees, each starting with an initial prompt message as the root node, which can have multiple child messages as replies, and these child messages can themselves have additional replies. All messages have a role property, either "assistant" or "prompter," with roles alternating strictly between "prompter" and "assistant" from prompt to leaf node.Data collection utilized a web app interface, dividing the process into five separate steps: prompting, labeling prompts, adding reply messages as prompter or assistant, labeling replies, and ranking assistant replies. To ensure high quality in the collected human responses, the OpenAssistant team provides guidelines for all contributors. These guidelines clarify the meanings, scales, and criteria for assigning labels and rankings during the labeling and ranking tasks, and stipulate that assistant responses should be polite, helpful, concise, friendly, and safety-aware.
@@
Stanfordnlp/SHP. SHP (Ethayarajh et al., 2022) is a dataset consisting of 385K collective human preferences regarding responses to questions and instructions across 18 subject areas, ranging from cooking to legal advice. These preferences are designed to reflect the helpfulness of one response compared to another and are intended for training RLHF reward models. Each entry includes a Reddit post featuring a question or instruction and a pair of top-level comments, where one comment is more favored by users. The dataset leverages the principle that if comment A is posted after comment B yet receives a higher score, A is generally considered more preferred. However, if A precedes B, we cannot make the same inference, as its higher score may simply result from increased visibility. SHP emphasizes preference labels that indicate which response is more helpful, contrasting with previous work that often focused on identifying less harmful responses.
@@
Tasksource/tasksource_dpo_pairs.The tasksource/tasksource_dpo_pairs dataset (Sileo, 2024) is specifically designed for use in DPO or RLHF, emphasizing expert-constructed examples over LLM-generated content. It encompasses data from various domains, particularly focusing on natural language inference (NLI) and logical reasoning tasks, to aid models in distinguishing between preferred and rejected responses. Structured in pairs, the dataset requires models to learn which of two responses is more favorable based on specific criteria. With over 4.8 million training examples, it also includes validation and test sets. Each example consists of a task description, a prompt, a chosen response, and a rejected response. This structure enables models to be fine-tuned to understand human-like preferences, enhancing their capability to generate useful and accurate responses in text generation tasks.
@@
Openbmb/RLHF-V-Dataset. RLHF-V-Dataset(Yu et al., 2024a) is a human-labeled dataset designed for the training of Multimodal Large Language Models(MLLMs). To address hallucination in MLLMs and enhance trustworthiness,  (Yu et al., 2024a) collect human-labeled fine-grained data for training and optimization, presenting RLHF-V. To be specific, RLHF-V collects high-quality fine-grained segment-level human corrections on the MLLMs responses with respect to question-answering and image description instructions. Human annotators are required to directly correct the hallucinated segments, which simultaneously introduces the preferred output. The final dataset contains a total of 5,733 preference pairs based on human annotation from 1.4k prompts from the instruction tuning dataset and image description from GPT-4. Using RLHF-V dataset in post-training effectively reduces hallucination in MLLMs.
@@
Anthropic Helpful and Harmless dialogue dataset(HH-RLHF). HH-RLHF(Bai et al., 2022a) is a human-labeled public dataset that contains 169,352 rows of human preference pair of dialogues between a human and an automated assistant, including human preference data collected in Training a Helpful and Harmless Assistant with Reinforcement Learning from Human Feedback (Bai et al., 2022a). For helpfulness, human crowdworkers are required to chat with base models to engage in dialogue task like answering a question and discussing interest topics. Then, crowdworkers need to choose the more helpful and honest response between the two generated responses. For the harmlessness or red-teaming dataset, crowdworkers are given instructions to elicit harmful responses from their models(similar to helpfulness but to choose harmful responses). The overall Anthropic HH-RLHF dataset also contains Human-generated and annotated red teaming dialogues from Red Teaming Language Models to Reduce Harms: Methods, Scaling Behaviors, and Lessons Learned (Ganguli et al., 2022).
@@ 4.2 AI labeled
Openbmb/RLAIF-V-Dataset. RLAIF-V-Dataset (Yu et al., 2024b) is a large-scale multimodal feedback dataset containing 83,132 high-quality preference pairs. The instructions are sourced from a diverse array of datasets, including MSCOCO (Lin et al., 2015), ShareGPT-4V (Chen et al., 2023a), MovieNet (Huang et al., 2020), Google Landmark v2 (Weyand et al., 2020), VQA v2 (Goyal et al., 2017), OKVQA (Marino et al., 2019), and TextVQA (Singh et al., 2019). Additionally, the dataset incorporates image description prompts introduced in RLHF-V as long-form image-captioning instructions. Training on this dataset enables multi-modal models to achieve greater trustworthiness compared to both open-source and proprietary models.
@@
Xinlai/Math-Step-DPO-10K. Math-Step-DPO-10K (Lai et al., 2024) is a high-quality step-wise AI-labeled preference dataset with 10795 rows of preference pairs that are tailored for mathematical reasoning. The dataset construction involves three main steps. Each data sample includes four entries: 1) prompt $x$; 2) initial reasoning steps; 3) preferred reasoning steps; and 4) undesirable reasoning steps. First, a mathematical problem prompt, accompanied by its ground-truth answer, is provided to a reference model, which generates a step-by-step solution using Chain-of-Thought (CoT). Next, solutions with steps differing from the ground-truth are selected to verify the correctness of each step, thereby identifying any erroneous steps. This process can be performed manually or with GPT-4. The undesirable reasoning steps are derived from these procedures. Multiple outputs are generated by inferring the reference model with the prompt $x$ and the preceding reasoning steps to obtain the corresponding correct reasoning step for each sample. This procedure can be repeated multiple times until the ground-truth answer is achieved.
@@
Openbmb/UltraFeedback. UltraFeedback (Cui et al., 2023a) is a large-scale, fine-grained, and diverse preference dataset designed for training robust reward and critic models.  Cui et al. (2023a) collected approximately 64,000 prompts from various sources, including UltraChat, ShareGPT, Evol-Instruct, TruthfulQA, FalseQA, and FLAN. These prompts were then used to query multiple LLMs, generating four different responses for each, resulting in a total of 256,000 samples.To ensure high-quality preference and textual feedback, they developed a fine-grained annotation instruction covering four aspects: instruction-following, truthfulness, honesty, and helpfulness. GPT-4 was employed to annotate the collected samples based on these guidelines. This dataset has been utilized by  Saeidi et al. (2024) for the chat completion task, comprising 63,000 pairs of selected and rejected responses for specific inputs. The UltraFeedback-binarized dataset is used to train alignment models employing the DPO algorithm.
@@
Argilla/Distilabel-capybara-dpo-7k-binarized. Capbara (Argilla, 2024) is a AI labelled dataset that contains 7536 conversations with chosen or rejected labels. Chosen and rejected pairs are formatted following OpenAI’s conversation format with potentially several turns between a user and an assistant.Multi-turn dialogue data is key to fine-tune capable chat models. Multi-turn preference data has been used by the most relevant RLHF works (Anthropic).However, there are very few multi-turn open datasets for DPO/RLHF. This dataset is the first of a series of datasets to design for DPO/RLHF fine-tuning. Capybara generates three responses to the last user message using OSS 7B models and distilabel’s LLMPool and the vLLM engine. From the 4 responses to each multi-turn dialogue, they use gpt-4-turbo to rank the quality of responses.
@@
Mlabonne/chatml-dpo-pairs. The dataset (Malabonne, 2024) is a pre-processed version of the Intel/orca-dpo-pairs dataset (Mukherjee et al., 2023), formatted using ChatML. It consists of 12,000 examples from the Orca-style dataset, which includes pairs of prompts and responses. The dataset is designed for use with Direct Preference Optimization (DPO), which utilizes a pair of responses with chosen and rejected labels. In this dataset, the data format is a system prompt, a user query, and two responses: one from ChatGPT (Bahrini et al., 2023) and one from Llama2-13b-chat (Touvron et al., 2023b). These responses are formatted, and the model is trained to prefer one response over the other, based on human or model feedback. The data has been processed to fit the ChatML format, making it suitable for various LLMs.
@@
Magpie-Align/Magpie-Llama-3.1-Pro-DPO-100K-v0.1. The Magpie-Llama-3.1-Pro-DPO-100K (Xu et al., 2024d) dataset contains 100,000 high-quality examples, making it suitable for fine-tuning LLMs using DPO, which involves comparing preferred and rejected responses to optimize the model’s behavior based on human-like preferences. The Magpie-Llama-3.1-Pro-DPO-100K-v0.1 dataset was constructed as part of the Magpie Alignment project, which focuses on synthesizing high-quality alignment data for training large language models (LLMs) without relying on manual prompt engineering or seed questions. Instead, this dataset is generated by prompting pre-aligned LLMs, such as Meta’s Llama 3.1 (AI@Meta, 2024), to autonomously produce instructional data.

### v3

@@ 4 Datasets
Datasets Task Description Modality Dataset Size (#Rows)
@@
Summarize from Human Feedback Stiennon et al. (2022) Summarization Text 193,841
OpenAI:Webgpt Nakano et al. (2022) Question Answering Text 19,578
Nvidia/HelpSteer Wang et al. (2023c) Preference Tuning Text 37,120
Nvidia/HelpSteer2 Wang et al. (2024b) Preference Tuning Text 21,362
Nvidia/HelpSteer3 Wang et al. (2025a) Preference Tuning Text 40,476
OpenAssistant/oasst1 Köpf et al. (2023) Conversation Text 88,838
Stanfordnlp/SHP Ethayarajh et al. (2022) Preference Tuning Text 385,563
Tasksource/Dpo-pairs Sileo (2024) Preference Tuning Text 5,128,939
RLHF-V-Dataset Yu et al. (2024a) Preference Tuning Multimodal 5,733
HH-RLHF Bai et al. (2022a) Preference Tuning Text 169,352
MM-RLHF Zhang et al. (2025) Preference Tuning Multimodal 16,342
Chatbot Arena Zheng et al. (2023a) Model Evaluation Text 33,000
Alpaca Farm Human Dubois et al. (2024) Instruction Tuning Text 101,000
MMSafe-PO Li et al. (2025a) Safety Preference Multimodal 5,667
@@
RLAIF-V-Dataset Yu et al. (2024b) Preference Tuning Multimodal 33,835
Math-Step-DPO Lai et al. (2024) Mathematical Reasoning Text 10,795
UltraFeedback Cui et al. (2023a) Preference Tuning Text 63,967
Distilabel-capybara-dpo Argilla (2024) Conversation Text 7,563
Chatml-dpo-pairs Mukherjee et al. (2023) Instruction Tuning Text 12,859
Magpie-Llama-3.1-DPO Xu et al. (2024d) Instruction Tuning Text 100,000
Tldr-preference-dpo davanstrien (2024) Summarization Text 522
Zsql-postgres-dpo Zerolink (2024a) Text-to-SQL Text 259,326
Truthy-dpo Zerolink (2024b) Truthfulness Preference Text 1,016
Py-dpo Jondurbin (2024a) Code Generation Text 9,466
MMInstruction/VLFeedback Li et al. (2023a) Instruction Tuning Multimodal 80,258
sqrti/SPA-VL Zhang et al. (2024a) Safety Preference Multimodal 100,788
Gutenberg-DPO Jondurbin (2024b) Creative Writing Text 918
Imdb-dpo Yuasosnin (2024) Summarization Text 10,000
RewardBench Lambert et al. (2024) Model Evaluation Text 8,108
RewardBench2 Malik et al. (2025) Model Evaluation Text 1,865
VL-RewardBench Li et al. (2025b) Model Evaluation Multimodal 1,247
Stack Exchange Preference Lambert et al. (2023) Instruction Tuning Text 10,741,532
Nectar Zhu et al. (2023) Preference Tuning Text 182,954
OrcaSlimOrca Lian et al. (2023) Instruction Tuning Text 517,982
Alpaca Farm GPT-4 Dubois et al. (2024) Instruction Tuning Text 20,000
CodeSteer-DPO-Dataset Chen et al. (2025) Code Generation Text 4,462
ANAH Gu et al. (2025) Hallucination Detection Text 783
Unified Preference Dataset Wang et al. (2025b) Preference Tuning Multimodal 236,000
MM-IFDPO-23k Ding et al. (2025) Instruction Tuning Multimodal 22,555
@@
DPO algorithm aligns LLMs using datasets that consist of preference pairs. Typically, such a dataset includes a prompt $x$ paired with two responses: a preferred (chosen) answer and a dispreferred (rejected) answer. Typically, human annotators are guided to evaluate and provide preferences among multiple responses generated from the same prompt. These datasets are denoted as human-labeled datasets in this survey. This process, however, is resource-intensive, especially when dealing with large volumes of data. To mitigate the high cost associated with human annotation, some datasets employ language models to synthesize preference pairs Cui et al. (2023a) Yu et al. (2024b) Zhang et al. (2024a). These synthesized datasets are referred to as AI-labeled datasets in this survey. In this section, we will explore the two types of datasets in greater detail.
@@ 4.1 Human labeled
Figure 11: Two examples of Human labeled datasets: (a) RLHF-V-Dataset is a Pair-Wise preference dataset, where the unmodified large model responses are marked as rejected items, and the responses that humans have finely tuned are marked as chosen items. (b) HelpSteer3 is a Point-Wise preference dataset, with responses generated by 17 models and annotated by over 6,400 people. A negative score indicates that R1 is better, while a positive score indicates that R2 is better. The absolute value of the score reflects the degree of superiority or inferiority.
@@
OpenAI:Summarize From Human Feedback. In downstream task like summarization, public online forum like Reddit can provide abundant text sources with a variety of topics. The TL;DR summarization dataset Völske et al. (2017) contains 3 millions post from reddit forum. To enhance the quality of data, Stiennon et al. Stiennon et al. (2022) filtered the dataset and left 123,169 posts with  5% as a validation set. Then  Stiennon et al. (2022) collect human feedback on the filtered TL;DR dataset. To improve human labelers’s responses on the instruction, Stiennon team answered human labeler’s questions in a shared chat room, and provided regular feedback on their performance, which ensures high agreement on the judgements of responses. The new collected human preference dataset contains 64k text summarization examples including human-written responses and human-rated model responses, which could server as high quality sources of training data for DPO.
@@
OpenAI:Webgpt comparisons. Webgpt comparisons Nakano et al. (2022) consist of 19,578 comparisons that are deemed appropriate for reward modeling, aiming at training a long-form question answering model aligned with human preferences. Each example in the dataset includes a pair of model-generated answers for a given question, along with associated metadata. Each answer is assigned a preference score from human annotators, allowing for the identification of the superior response. To gather these human preferences,  Nakano et al. (2022) hire well-educated annotators, typically holding at least an undergraduate degree, due to the complexity of the tasks involved. Ultimately, this dataset captures human preferences in the generation of WebGPT, facilitating the optimization of the question-answering model to better align with human expectations.
@@
Nvidia/HelpSteer. HelpSteer Wang et al. (2023c) is an open-source Helpfulness Dataset (CC-BY-4.0) designed to enhance model alignment for improved helpfulness, factual accuracy, and coherence, while allowing for adjustments in response complexity and verbosity. The dataset comprises 37,120 samples, each containing a prompt, a generated response, and five human-annotated attributes rated from 0 to 4, with higher scores indicating better performance.Prompts are sourced from various origins, with approximately half generated by Scale AI. The remaining prompts come from Open Question Answering, Generation, and Brainstorming tasks, as well as five specific tasks: Rewrite, Summarization, Classification, Extraction, and Closed Question Answering. Responses are produced by an early version of a 43 billion parameter in-house model, which generates up to four responses per prompt using sampling techniques to ensure diversity while maintaining reasonableness. The annotation process involved around 200 U.S.-based human annotators, who evaluated each response based on five attributes: Helpfulness, Correctness, Coherence, Complexity, and Verbosity, using a Likert-5 scale from 0 to 4.
@@
Nvidia/HelpSteer2. HelpSteer2 Wang et al. (2024b), an open-source helpfulness dataset designed to train state-of-the-art reward models, comprises 21,362 samples, each containing a prompt, a response, and five human-annotated attributes, rated from 0 to 4, with higher scores indicating better performance. Consecutive samples (e.g., sample 1 with sample 2, sample 3 with sample 4) share the same prompt, allowing for preference pairs based on the helpfulness score (e.g., for training DPO or Preference RM), in addition to training SteerLM Regression RM. Over 95% of the prompts are sourced from ShareGPT RyokoAI (2023) , a platform where ChatGPT users voluntarily share their conversations, while a small proportion ( 5%) is generated by humans from Scale AI.Responses are produced by early versions of a mix of ten different in-house LLMs, generating two responses per prompt (one from each model) using sampling techniques to provide diverse yet reasonable answers.
Nvidia/HelpSteer3. HelpSteer3 Wang et al. (2025a), contains 40,476 preference samples, each containing a domain, language, context, two responses, an overall preference score between the responses, as well as individual preferences from up to 3 annotators. Each individual preference contains a preference score in addition to a concise reasoning for their preference in 1-2 sentences. 95% of the data is the training set and 5% is the validation set. The prompts are sourced from ShareGPT and WildChat-1M, covering four aspects: General tasks, Code, Multilingual, and STEM (Science, Technology, Engineering, and Mathematics). Responses were generated from a variety of 17 commercially-permissive models.
@@
OpenAssistant/oasst1. OpenAssistant Conversations (OASST1) Köpf et al. (2023) is a human-generated, human-annotated assistant-style conversation corpus comprises 161,443 messages in 35 different languages, annotated with 461,292 quality ratings, resulting in over 10,000 fully annotated conversation trees. The corpus is the product of a worldwide crowd-sourcing initiative involving over 13,500 volunteers.The dataset consists of message trees, each starting with an initial prompt message as the root node, which can have multiple child messages as replies, and these child messages can themselves have additional replies. All messages have a role property, either "assistant" or "prompter," with roles alternating strictly between "prompter" and "assistant" from prompt to leaf node.Data collection utilized a web app interface, dividing the process into five separate steps: prompting, labeling prompts, adding reply messages as prompter or assistant, labeling replies, and ranking assistant replies. To ensure high quality in the collected human responses, the OpenAssistant team provides guidelines for all contributors. These guidelines clarify the meanings, scales, and criteria for assigning labels and rankings during the labeling and ranking tasks, and stipulate that assistant responses should be polite, helpful, concise, friendly, and safety-aware.
@@
Stanfordnlp/SHP. SHP Ethayarajh et al. (2022) is a dataset consisting of 385K collective human preferences regarding responses to questions and instructions across 18 subject areas, ranging from cooking to legal advice. These preferences are designed to reflect the helpfulness of one response compared to another and are intended for training RLHF reward models. Each entry includes a Reddit post featuring a question or instruction and a pair of top-level comments, where one comment is more favored by users. The dataset leverages the principle that if comment A is posted after comment B yet receives a higher score, A is generally considered more preferred. However, if A precedes B, we cannot make the same inference, as its higher score may simply result from increased visibility. SHP emphasizes preference labels that indicate which response is more helpful, contrasting with previous work that often focused on identifying less harmful responses.
@@
Tasksource/tasksource_dpo_pairs.The tasksource/tasksource_dpo_pairs dataset Sileo (2024) is specifically designed for use in DPO or RLHF, emphasizing expert-constructed examples over LLM-generated content. It encompasses data from various domains, particularly focusing on natural language inference (NLI) and logical reasoning tasks, to aid models in distinguishing between preferred and rejected responses. Structured in pairs, the dataset requires models to learn which of two responses is more favorable based on specific criteria. With over 4.8 million training examples, it also includes validation and test sets. Each example consists of a task description, a prompt, a chosen response, and a rejected response. This structure enables models to be fine-tuned to understand human-like preferences, enhancing their capability to generate useful and accurate responses in text generation tasks.
@@
Openbmb/RLHF-V-Dataset. RLHF-V-DatasetYu et al. (2024a) is a human-labeled dataset designed for the training of Multimodal Large Language Models(MLLMs). To address hallucination in MLLMs and enhance trustworthiness,  Yu et al. (2024a) collect human-labeled fine-grained data for training and optimization, presenting RLHF-V. To be specific, RLHF-V collects high-quality fine-grained segment-level human corrections on the MLLMs responses with respect to question-answering and image description instructions. Human annotators are required to directly correct the hallucinated segments, which simultaneously introduces the preferred output. The final dataset contains a total of 5,733 preference pairs based on human annotation from 1.4k prompts from the instruction tuning dataset and image description from GPT-4. Using RLHF-V dataset in post-training effectively reduces hallucination in MLLMs.
@@
Anthropic Helpful and Harmless dialogue dataset(HH-RLHF). HH-RLHFBai et al. (2022a) is a human-labeled public dataset that contains 169,352 rows of human preference pair of dialogues between a human and an automated assistant, including human preference data collected in Training a Helpful and Harmless Assistant with Reinforcement Learning from Human Feedback Bai et al. (2022a). For helpfulness, human crowdworkers are required to chat with base models to engage in dialogue task like answering a question and discussing interest topics. Then, crowdworkers need to choose the more helpful and honest response between the two generated responses. For the harmlessness or red-teaming dataset, crowdworkers are given instructions to elicit harmful responses from their models(similar to helpfulness but to choose harmful responses). The overall Anthropic HH-RLHF dataset also contains Human-generated and annotated red teaming dialogues from Red Teaming Language Models to Reduce Harms: Methods, Scaling Behaviors, and Lessons Learned Ganguli et al. (2022).
MM-RLHF. MM-RLHF Zhang et al. (2025) is a dataset containing 120k fine-grained, human-annotated preferences. It aims to align multimodal large language models with human preferences. This dataset was created by clustering and sampling from three areas (image understanding, video understanding, and multimodal safety), resulting in 30,000 high-quality queries. Additionally, MM-RLHF employs the most advanced response generation models to ensure response quality. Finally, after annotation, scoring, and ranking by over 50 professional annotators, the final dataset was obtained. Notably, this dataset provides the reasons for all scores and rankings, enhancing interpretability.
MMSafe-PO. Li et al. (2025a) MMSafe-PO comprises 5,667 multimodal instructions, each containing a text and a corresponding image. For each instruction, there is a chosen response and a rejected response.This dataset is transformed from the Anthropic-HH dataset, which recognizes the entities in text instructions and generates images of them, making this dataset applicable to the training of MLLMS. Like the Anthropic-HH dataset, it takes into account quantity, quality, and the harmless alignment objective.
@@ 4.2 AI labeled
Figure 12: Two examples of AI labeled datasets: (a) RLAIF-V-Dataset is a Pair-Wise preference dataset. It designed a strategy to decompose responses, first breaking down responses into atomic claims and converting them into polar questions. This way, large models only need to reply with yes or no for score calculation, and ultimately form preference pairs based on the score’s ranking. (b) ULTRAFEEDBACK is a Point-Wise preference dataset. It generates responses using multiple powerful large models and uses detailed prompts to have GPT-4 rank and score each group of responses on four dimensions to form the final preference set.
@@
Openbmb/RLAIF-V-Dataset. RLAIF-V-Dataset Yu et al. (2024b) is a large-scale multimodal feedback dataset containing 83,132 high-quality preference pairs. The instructions are sourced from a diverse array of datasets, including MSCOCO Lin et al. (2015), ShareGPT-4V Chen et al. (2023a), MovieNet Huang et al. (2020), Google Landmark v2 Weyand et al. (2020), VQA v2 Goyal et al. (2017), OKVQA Marino et al. (2019), and TextVQA Singh et al. (2019). Additionally, the dataset incorporates image description prompts introduced in RLHF-V as long-form image-captioning instructions. Training on this dataset enables multi-modal models to achieve greater trustworthiness compared to both open-source and proprietary models.
@@
Xinlai/Math-Step-DPO-10K. Math-Step-DPO-10K Lai et al. (2024) is a high-quality step-wise AI-labeled preference dataset with 10795 rows of preference pairs that are tailored for mathematical reasoning. The dataset construction involves three main steps. Each data sample includes four entries: 1) prompt $x$; 2) initial reasoning steps; 3) preferred reasoning steps; and 4) undesirable reasoning steps. First, a mathematical problem prompt, accompanied by its ground-truth answer, is provided to a reference model, which generates a step-by-step solution using Chain-of-Thought (CoT). Next, solutions with steps differing from the ground-truth are selected to verify the correctness of each step, thereby identifying any erroneous steps. This process can be performed manually or with GPT-4. The undesirable reasoning steps are derived from these procedures. Multiple outputs are generated by inferring the reference model with the prompt $x$ and the preceding reasoning steps to obtain the corresponding correct reasoning step for each sample. This procedure can be repeated multiple times until the ground-truth answer is achieved.
@@
Openbmb/UltraFeedback. UltraFeedback Cui et al. (2023a) is a large-scale, fine-grained, and diverse preference dataset designed for training robust reward and critic models.  Cui et al. (2023a) collected approximately 64,000 prompts from various sources, including UltraChat, ShareGPT, Evol-Instruct, TruthfulQA, FalseQA, and FLAN. These prompts were then used to query multiple LLMs, generating four different responses for each, resulting in a total of 256,000 samples.To ensure high-quality preference and textual feedback, they developed a fine-grained annotation instruction covering four aspects: instruction-following, truthfulness, honesty, and helpfulness. GPT-4 was employed to annotate the collected samples based on these guidelines. This dataset has been utilized by  Saeidi et al. (2024) for the chat completion task, comprising 63,000 pairs of selected and rejected responses for specific inputs. The UltraFeedback-binarized dataset is used to train alignment models employing the DPO algorithm.
@@
Argilla/Distilabel-capybara-dpo-7k-binarized. Capbara Argilla (2024) is a AI labelled dataset that contains 7536 conversations with chosen or rejected labels. Chosen and rejected pairs are formatted following OpenAI’s conversation format with potentially several turns between a user and an assistant.Multi-turn dialogue data is key to fine-tune capable chat models. Multi-turn preference data has been used by the most relevant RLHF works (Anthropic).However, there are very few multi-turn open datasets for DPO/RLHF. This dataset is the first of a series of datasets to design for DPO/RLHF fine-tuning. Capybara generates three responses to the last user message using OSS 7B models and distilabel’s LLMPool and the vLLM engine. From the 4 responses to each multi-turn dialogue, they use gpt-4-turbo to rank the quality of responses.
@@
Mlabonne/chatml-dpo-pairs. The dataset Malabonne (2024) is a pre-processed version of the Intel/orca-dpo-pairs dataset Mukherjee et al. (2023), formatted using ChatML. It consists of 12,000 examples from the Orca-style dataset, which includes pairs of prompts and responses. The dataset is designed for use with Direct Preference Optimization (DPO), which utilizes a pair of responses with chosen and rejected labels. In this dataset, the data format is a system prompt, a user query, and two responses: one from ChatGPT Bahrini et al. (2023) and one from Llama2-13b-chat Touvron et al. (2023b). These responses are formatted, and the model is trained to prefer one response over the other, based on human or model feedback. The data has been processed to fit the ChatML format, making it suitable for various LLMs.
@@
Magpie-Align/Magpie-Llama-3.1-Pro-DPO-100K-v0.1. The Magpie-Llama-3.1-Pro-DPO-100K Xu et al. (2024d) dataset contains 100,000 high-quality examples, making it suitable for fine-tuning LLMs using DPO, which involves comparing preferred and rejected responses to optimize the model’s behavior based on human-like preferences. The Magpie-Llama-3.1-Pro-DPO-100K-v0.1 dataset was constructed as part of the Magpie Alignment project, which focuses on synthesizing high-quality alignment data for training large language models (LLMs) without relying on manual prompt engineering or seed questions. Instead, this dataset is generated by prompting pre-aligned LLMs, such as Meta’s Llama 3.1 AI@Meta (2024), to autonomously produce instructional data.